OpenStudy (anonymous):

In a circle of radius 8, the length of the arc that subtends a central angle of 183 degrees is

OpenStudy (anonymous):

I believe the formula is \[s = r \theta\] Where theta is in radians.. Can anyone confirm this??

OpenStudy (anonymous):

Convert 183 degrees into radians \[183degrees = \frac{183\pi}{180}radians\] so the arc length s would equal \[s = 8(\frac{183\pi}{180})\] So s would equal \[s = 25.55\]

OpenStudy (anonymous):

Thats what i got but im not sure if its right.

OpenStudy (anonymous):

yeah it is actually A=(theta/360)x (2) (pi) (r)

OpenStudy (anonymous):

Ok i did it right then.. yayy lol

OpenStudy (anonymous):

yeah I appologize Im not good at using the equation thing here so it takes a minute for me to type it out

OpenStudy (anonymous):